SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 1 Jun 2004

Status Report From: NASA HQ
Posted: Tuesday, June 1, 2004

image

All ISS systems continue to function nominally except those noted previously or below.   Underway: Week 5 of Expedition 9.

As part of station inspection in the morning, CDR Gennady Padalka did the regular (monthly) routine checkup in the DC-1 "Pirs" docking module on the AZS circuit breakers on the BVP Amp Switch Panel -- they should all be On -- and the LEDs (light-emitting diodes) of the 14 fuses in Fuse Panels BPP-30 & BPP-36 (last time done: 5/1).

After station inspection, before breakfast and first exercise, both crewmembers performed their third session of the periodic Russian biomedical assessments PZEh-MO-8 (body mass measurement) and PZEh-MO-7 (calf volume measurement).  [Padalka set up the BMM mass-measuring device, which uses calibrated springs to determine the subject's mass in weightless space, and stowed it away after the tests.  Calf measurements (left leg only) are taken with the ISOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference points, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures.]

The crew again spent several hours with cargo transfer operations from Progress 14P, tracking the movements with the IMS (inventory management system) and referring to a detailed unpack & stowage list uplinked from TsUP/Moscow.   [The unpack list is supplemented by an U.S. list specifying transfers & stowage of CHeCS IMAK (Crew Health Care Systems/ISS Medical Accessory Kit) items intended for the AMP (Ambulatory Medical Pack) and ALSP (Advanced Life Support Pack).]

In support of subsequent payload operations with EXPRESS rack 5 (ER5), Fincke powered up the ER5 laptop and assisted with the rack's remote activation from POIC (Payload Operations & Integration Center)/Huntsville.  For the next 10 hrs, the ground tracked the power and thermal resources of the EXPRESS rack during its operations.  This included a run of the SNFM (Serial Network Flow Monitor) for which the Science Officer activated the SNFM on the ER5 laptop computer (ELC) for a 3-hr. untended autocapture of LAN-1 science network "bridged" data traffic during the day.  Mike then powered ER5 down again.

The CDR completed the periodic replenishing of the Elektron's water supply for electrolysis, filling the KOV thermal loops' EDV container with purified (deionized) water from the BKO multifiltration/purification column unit.  (Last time done: 5/18)   [The procedure is specifically designed to prevent air bubbles from getting into the BZh liquid unit where they could cause micropump impeller cavitation and Elektron shutdown, as in the past on numerous times.  In the procedure, the EDV water is drawn from the BKO and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number).]

After reconfiguring an SSC (station support computer) in the Airlock for EMU (extravehicular mobility unit) battery operations, Mike Fincke started a discharge on three selected PGT (pistol grip tool) batteries for the upcoming Orlan spacewalk.   [The selected NiMH (nickel metal hydride) batteries are ##1005, 1015, and 1018.  Last Friday's (5/28) checkout of Battery Charger 2 (BC2) determined that there is no issue with BC2 itself, clearing it for today's initiation of the maintenance discharge and the subsequent recharge starting on 6/3.]

As is standard practice after Russian dockings, CDR Padalka took two photos of the Service Module (SM) aft-end docking assembly (StA) used for the Progress M-49/14P linkup.  These images will be used to refine current understanding of docking conditions.   [The objective was to take photo imagery of the scratch or scuff mark left by the head of the docking probe on the internal surface of the drogue (docking cone) ring.  As other cosmonauts before him, Gennady used the Kodak 760 digital still camera to take two pictures with the hatch closed down and downlinked them later via OCA.]

The FE performed the daily leak check on the Lab window pressure test equipment.  Today's reading was 9.4 Torr (mmHg) or 0.18 psi.

Padalka terminated the current measurement session of the Molniya-SM/LSO equipment, running on SM window #3 since 5/28, and deactivated its hardware, powering off the EGE-1 laptop.  Copying of data onto disk for return is scheduled for tomorrow.   [Purpose of Molniya-SM, similar to the French LSO experiment, is to record storm phenomena and other related events in the Earth's equatorial regions.  The experiment requires orbital-sighting predictions using an up-to-date NORAD tracking TLE (two-line element) provided by NASA.  Objective of LSO was to study rare optical phenomena occurring in the upper layers of Earth's atmosphere, so-called "sprites" (i.e., puzzling glow phenomena observed above thunderstorm clouds).]

Gennady also completed the daily routine maintenance of the SM's SOZh life support system, while Mike prepared the daily "delta" file for updating the IMS (inventory management system) database(s).

The crew worked out on TVIS treadmill and RED exerciser, in accordance with their daily 2.5-hr. physical exercise protocol.

At 5:30am EDT, ISS attitude control was handed over to the Russian motion control system in the SM for the standard dynamic test of the US-21 matching unit installed in the Progress 14P on 5/28.  [The test firings of Progress thruster manifolds 1 and 2 were completed at 9:13am, and control authority handed back to U.S. CMGs (control moment thrusters) at 9:30am to resume momentum management in LVLH (local vertical/local horizontal).  Progress thruster yaw and pitch control was incorporated into the SM's propulsion control algorithm.]

Overnight the ground uplinked an updated list of 21 "yellow tag" items, including their constraints on usage aboard ISS, for an on-orbit review by the crew.  Before a yellow-tagged item can be used, the ground needs to be consulted.   [Yellow tags, more formally called "uncertified dual ops tags", are used to identify items not certified for ISS Operations (certification and/or paperwork not complete prior to launch); items which have IP (international partner) segment-specific certification (can be used in one IP segment but should not be used in anther IP segment); items that could pose a safety hazard; and items that are broken or expired.  Blank yellow tags are flown so hardware can be tagged on-orbit as necessary.]

CEO (Crew Earth Observations): no targets for today due to yesterday's Memorial Day holiday.

U.S. and Russian Segment Status  (as of today, 12:50pm EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) -- 742; temperature (deg C) -- 25.2; ppO2 (mmHg) -- 158.9; ppCO2 (mmHg) -- 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) -- 751; temperature (deg C) -- 20.2.
  • FGB Cabin:  Pressure (mmHg) -- 756; temperature (deg C) -- 21.7.
  • Node:  Pressure (mmHg) -- 751.52; temperature (deg C) -- 22.6 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 753.4; temperature (deg C) -- 24.9; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) -- 753.53; temperature (deg C) -- 21.8; shell heater temp (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) -- n/a
  • PMA-2:  Shell heater temp (deg C) -- n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed position ("blind", non solar-tracking dual-angle mode with "night glider"/"sun slicer" 60 deg. drag reduction sweep).
  • SM batteries:  All batteries (8) are on line in "Partial Charge" mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3436 kg (7575 lb) as of 5/20/04;  [SM(552) + FGB(2884) + Progress M-1(0)].  (Capability: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:12am EDT [= epoch]):

  • Mean altitude -- 362.8 km
  • Apogee -- 366.7 km
  • Perigee -- 358.8 km
  • Period -- 91.8 min.
  • Inclination (to Equator) --  51.6336 deg
  • Eccentricity -- 0.0005869
  • Solar Beta Angle --  -2.5 deg
  • Orbits per 24-hr. day -- 15.69
  • Mean altitude loss in last 24 hours -- 100 m
  • Revolutions since FGB/Zarya launch (Nov. '98)  -- 31584

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see  http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks